# revtex4-2Repair the float

### v1

@@ II.8 Principled Embedding Dimension via Information Criteria
where $K$ is the number of parameters, $n:=\binom{N}{2}$ the number of observations and $\mathcal{L}$ the likelihood of the model. They encode the trade-off between the goodness-of-fit $-\ln(\mathcal{L})$ and the complexity of the model. Therefore, the “best model” is the one with the minimum AIC or BIC; which one of the two remains a debated choice: the AIC is asymptotically equivalent to the Kullback-Leibler divergence among the generating model and a candidate one [27] whereas the BIC to the Description Length (DL) [28]. As the DL is the only one embodying the “trade-off” paradigm, we decided to select the minimum BIC criterion. In Equation 38, the scores are not comparable across scale, therefore, we normalized the AIC and BIC scores as
@@
II.9 Statistical Impossibility into the application of Train and Test Split
Since every pair is seen as a Bernoulli random variable not identically distributed [41, 8], this implies that every link has to be accounted in the training procedure since it is not already modelled by other pairs.
III World Trade Web results
@@
Figure 3: Left: Visualization of the Equation 13 where the LHS lies on the y-axis and the RHS on the x-axis. Right: Cross comparison among LPCA-(8,8) and MSM-16 focusing on the CC. The upper panel reports the expected $\left\{\langle CC_{i}\rangle\right\}_{j\in[1,N]}$ at level 0 whereas the lower at level 4. In addition, the first row refers to the LPCA-(8,8) and the second to MSM-16. The inset highlights the $\widetilde{\mathbf{P}}^{(\ell)}$ against $\hat{\mathbf{P}}^{(\ell)}$ at that resolution level, namely 00 or $4$.
@@
Figure 4: Fundamental Network Measurements at level 4 for the ION dataset. In the upper panel, the x-axis hosts the observed measurements whereas the y-axis the expected ones. In particular, from the left one finds the DEG, the ANND and the CC. The single inset depicts the scattered plots of $\widetilde{\mathbf{P}}^{(4)}$ against $\hat{\mathbf{P}}^{(4)}$. In the lower panel, it is displayed the behavior of the network measurements as the degrees increase either for the observed quantities and the expected ones. One finds every expected value calculated with the fitted $\hat{\mathbf{P}}^{(4)}$ (blue) and the summed $\widetilde{\mathbf{P}}^{(4)}$ (orange) while the observed measures are depicted in azure. The left half of the figure refers to the LPCA-(8,8) model, whereas the right half of the figure to the MSM-16.
@@
Figure 5: Confusion matrices, ROC and PR curves at level 4 for the ION dataset. The left most side is occupied by the two confusion matrices for the fitted $\hat{\mathbf{P}}^{(4)}$ (upper) and the summed $\widetilde{\mathbf{P}}^{(4)}$. The middle plot (ROC curves) reports the behavior of the TPR VS FPR as the threshold goes from 1 to 0 (reading the graphic left-to-right). As before, the two curves are associated with the fitted (blue) model and the summed (orange) one. The left half of the figure refers to the LPCA-(8,8) model, while the right half of the figure to the MSM-16.
@@
Figure 6: Left: Reconstruction Accuracy (y-axis) by model, level and network statistics. Right: Area Under the ROC and PR curves for the summed models by diminishing the number of nodes or, equivalently, increasing the scale. More concretely, the third point reports the AUC-ROC and PR from Figure 5 as y-coordinates while the x-coordinates are the number of nodes $N_{4}$.
@@
Figure 7: Left: Rich-Club-Triangle Density at level 0. The plot shows the evolution of the $\rho^{(\ell)}(c)$ with respect to the degree $k$ in azure solid dots. The other markers identify the expected $\langle\rho^{(\ell)}(c)\rangle$ by the models: LPCA-(1,1) (azure circles), LPCA-(8,8) (orange triangles), MSM-1 (green circles), MSM-4 (red circles), MSM-8 (violet crosses), MSM-16 (brown triangles). Right: Rich-Club-Triangle Density at level 4.

### v2

@@ Appendix D Principled Embedding Dimension via Information Criteria
where $K$ is the number of parameters, $n:=\binom{N}{2}$ the number of observations and $\mathcal{L}$ the likelihood of the model. They encode the trade-off between the goodness-of-fit $\ln(\mathcal{L})$ and the complexity of the model. Therefore, the “best model” is the one with the minimum AIC or BIC; which one of the two remains a debated choice: the AIC is asymptotically equivalent to the Kullback-Leibler divergence among the generating model and a candidate one [24] whereas the BIC to the Description Length (DL) [25]. As the DL is the only one embodying the “trade-off” paradigm, we decided to select the minimum BIC criterion. In Equation 87, the scores are not comparable across scale, therefore, we normalized the AIC and BIC scores as
@@
Appendix E World Trade Web results
@@
Figure 10: Upper Plot: This figure illustrates the numerical evaluation of the two sides of the scale-invariant equation at level $\ell=2$ for the $LPCA$ and MSMMSM models. More precisely, the left-hand side (LHS) is represented on the y-axis while the right-hand side (RHS) on the x-axis. Lower Plot: Cross comparison of LPCA-(8,8) and MSM-16 in predicting the clustering coefficient (CC) (see the main text). The upper panel reports the expected clustering coefficient at level 0, while the lower panel depicts its corresponding values at level 2. The first column refers to the LPCA-(8,8), whereas the second to the MSM-16.
@@
Figure 11: The plots display the predicted network measurements for the WTW dataset at level 4 according to LPCA-(8,8) (left plots) and the MSM-16 (right plot). In each of the upper panels, the x-axis represents the observed measurements, whereas the y-axis shows the expected ones. From left to right, the plots illustrates the degree (DEG), the average-nearest-neighbor degree (ANND), and the clustering coefficient ($C$). The scatter plot comparing $\mathbf{P}_{sum}^{(2)}$ with $\hat{\mathbf{P}}^{(2)}$ is reported in the inset. The lower panels show the behavior of the network measurements as the degrees increase. The observed values are colored in blue, while the one calculated using the fitted $\hat{\mathbf{P}}^{(2)}$ model in red or the summed $\mathbf{P}_{sum}^{(2)}$ model in orange. Additionally, the z-score of the predicted number of links is indicated in the legend of the upper-left plot.
@@
Figure 12: The graphs show the Confusion matrices, ROC and PR curves for the ION dataset at level 2 obtained by means of the LPCA-(8,8) (upper plots) and the MSM-16 (lower plot). On the left most side, one may display the two confusion matrices for the fitted $\hat{\mathbf{P}}^{(2)}$ (upper) and the summed $\mathbf{P}_{sum}^{(2)}$. The middle plot reports the Receiver-Operator Curve, while the right-most plot depicts the Precision Recall curve. As in the previous plot, the two curves are associated with the fitted (red) model and the summed (orange) one.
@@
Figure 13: Left: Reconstruction Accuracy (y-axis) by model, level and network statistics. Right: Area Under the ROC and PR curves (y-axis) for the summed models as the scales increase (x-axis) or, equivalently, the numbers of nodes diminish.
@@
Figure 14: Left: Triangle Density at level 0. The plot shows the evolution of the $\tau^{(\ell)}(c)$ with respect to the degree $c$ in blue solid dots. The other markers identify the expected $\langle\tau^{(\ell)}(c)\rangle$ by the LPCA-(1,1), LPCA-(8,8), MSM-1, MSM-2, MSM-8, MSM-16. Right: Triangle Density at level 2.